# Reference Points, Risk-Taking Behavior, and Competitive Outcomes in Sequential Settings

### v3

@@ 2.1 Data source
We use the data from OpenPowerlifting which is a community service project to create a permanent and open archive of the world’s powerlifting data. [^5] https://gitlab.com/openpowerlifting. The OpenPowerlifting database includes a wide range of data fields of official powerlifting competition. It includes competition details such as federation name, date and location, level (local, national, international), athlete information such as name, age, weight class, gender, lift results of squat, bench press, and deadlift (with attempted weights and success/failure status), Wilks score and other relative strength metrics, and rankings of placement within the competition historical rankings and records.
In this paper, we focus on the bench press competition which is the most popular division and less complex than Squat-Bench-Deadlift (SBD) composite competition. We use all official attempts data of 15-69 aged lifters at competition categories based on age-class, weight-class, and gender, in which the number of participant lifters is at least two to capture competition pressure.
@@
Sources: The OpenPowerlifting database. We also observe Wilks score and other relative strength metrics, and rankings of placement within the competition historical rankings and records used for robustness check.
@@
Personal best weights of lifters who have never participated in an official bench press competition are recorded as zero. In the empirical analysis, we focus on the second and third attempts and treat the realized outcome in the first attempt as the personal best for such lifters because the success probability in the first attempt is 86% for raw lifters, 77% for single-ply lifters, and 70% for multi-ply lifters.
@@ 2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve their own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition. [^6] See details in the Internatinal Powerlifting Federation (IPF) Technical Rules Book.
@@
Each lifter has three attempts.777 Lifting Procedure for each lifter at each attempt is officially divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all the rules and completes the lift as per the judge’s commands from the viewpoints of two of three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
@@
Notes: The box means that the player appears on the stage. $W_{it}^{j}$ is lifter $i$’s attempt weight in $j$-th attempt in competition $t$.
To highlight the available information of each lifter at each attempt, consider a competition with three lifters (A, B, and C). The timeline of the first and second attempts is shown in Figure 1. For opening attempts, lifters $A,B$, and $C$ declare 100 kg, 105 kg, and 110 kg before the competition. Then, the order will be $A,B$, and $C$. Lifter $A$ attempts 100 kg then the outcome is realized. Next, lifter $B$ attempts 105 kg. At the same time, lifter $A$ declares the next attempt weight 110 kg in one minute conditional on $A$’s own first outcome and $B$ and $C$’s attempt weights before $B$’s outcome is realized because lifter $B$ is trying 105 kg at that time. Lifter $B$’s outcome is realized. Next, lifter $C$ attempts 110 kg. At the same time, lifter $B$ declares the next attempt weight 115 kg in one minute knowing $B$’s own first outcome, $A$’s outcome and next attempt, and $C$’s attempt weight before $C$’s outcome is realized because lifter $C$ is trying 110 kg at that time. Lifter $C$’s outcome is realized. Lifter $C$ declares the next attempt weight 112.5 kg in one minute knowing $A$ and $B$s’ outcomes, their next attempts, and $C$’s own first outcome. For the second attempts, the new order will be A (110 kg), C (112.5 kg), B (115 kg). After the same procedure, the new order for the third attempts, will be similarly determined by the declared weights at the second attempts.
@@ 2.3 Exogenous pressure
The above sequential game structure in which each player does not know the subsequent attempts and outcomes provides exogenous variations to the choice of the attempt weight and the success of lifting the attempt weight, which is analogous to the lottery choice and realized gain of the lottery. We focus on the moves of the closest rivals to each lifter, that is, his lower rank and higher rank rival lifters, which are critical to finalize his rank in the competition.
@@
In the choice of attempt weight at the second attempt, first, each lifter knows the first attempt outcome and the second attempt weight of his one lower rank rival. These are uncorrelated with the lifter’s unobservable characteristics, such as his body condition on the competition date, because the rival declares his second attempt weight before observing the outcome of the lifter’s second attempt, which—if observable—could serve as a proxy for the lifter’s condition. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. Second, each lifter expects the second attempt weight of his one higher rank rival, although this value is not yet observed. We assume that each lifter can make a prediction of a higher-ranked rival’s action based on information about the rival’s past attempts and a rich set of covariates. [^8] The assumption is supported from our data in Section 7.2. Strategic reasoning in intermediate weight selections is also inherently limited, as each subsequent attempt must exceed the current attempt. While incorporating dynamic and strategic considerations is theoretically appealing, doing so introduces significant complexity and falls beyond the scope of our empirical analysis. Because these predictions are based solely on observable characteristics, they are independent of the lifter’s own unobserved factors, such as temporary performance conditions. As the first attempt weight of his one higher rank rival is heavier than his current best weight, it is more difficult to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain.
@@
In the second attempt of lifting the attempt weight, each lifter knows the first and second attempt outcomes of his one lower rank rival. These are uncorrelated with the lifter’s unobserved characteristics, such as his body condition on the competition date, because the lower-ranked rival completes both attempts without observing the lifter’s second outcome. The exogenous pressure affects the risk of being turned around by one lower rank rival, that is, the risk of loss. If his second attempt fails, he is turned around by the rival. Second, each lifter also knows the second attempt weight of his one higher rank rival. This declared weight is uncorrelated with the lifter’s unobserved characteristics, because at the time of declaration, the higher-ranked rival has not yet observed the lifter’s second outcome. If his second attempt is successful, he has more chances to turn around the rival. The exogenous pressure affects the risk of challenging turning around his one higher rank rival, that is, the risk of gain. Note that the pressure matters only in the lifting stage.
@@ 2.4 Data pattern
We illustrate data patterns of the choice of attempt weight and the success probability of the attempt weight conditional on the above exogenous pressure. For illustration, we show the case of the most popular competition category (Male, Raw, 24-39 age class). The data pattern is prevalent in the other competition categories. We omit the data pattern for the first bench press attempts because the success probability at that stage is more than 75%.
@@
In summary, the data suggests that when lifters face pressure in critical situations—such as attempting to improve their rank—they tend to choose more challenging weights. However, unlike the second attempt, where pressured lifters often maintain higher success rates, the third attempt shows a more substantial decline in overall success probabilities, suggesting increased difficulty in securing successful lifts under pressure.
@@
We examine the pressure exerted by rivals during competition. A natural question arises: who constitutes a rival for each lifter within a competition? Table 3 illustrates the distribution of rank changes for lifters between their second and third attempts. The majority of lifters maintain their rank, as shown by the high proportion of zero rank changes (67% in the second attempt and 72% in the third attempt). A smaller percentage of lifters experience a rank increase ($+1$) between attempts (14% in the second and 10% in the third), while the proportion of those gaining two or more ranks is notably lower (5% in the second and 3% in the third). Similarly, lifters dropping one rank ($-1$) account for 8% in the second attempt and 9% in the third, with even fewer dropping two or more ranks (5% and 6%, respectively). Based on this distribution, we focus on the immediate competitors, specifically the lifter ranked directly below and the one ranked directly above, as these rivals most significantly influence rank changes.
@@
3 Theoretical prediction
In this section, we derive empirical predictions based on expected utility theory and prospect theory. [^9] See the survey of O’Donoghue and Sprenger (2018) and DellaVigna (2018) for reference. The choice of an attempt weight is analogous to a binary lottery choice drawn from a set of lotteries $\left(w,q(w)\right)$, where $w$ represents the attempt weight and $q(w)\in[0,1]$ denotes the probability of success. We assume that the success probability is decreasing in weight, i.e., $q^{\prime}(w)<0$. Following Tversky and Kahneman (1992), we adopt a rank-dependent value function $v(\cdot)$ with reference point $r$ defined as
$v(x)=\begin{cases}(x-r)^{\alpha}&\mbox{if\ }x\geq r\\ -\lambda(r-x)^{\alpha}&\mbox{if\ }x<r,\end{cases}$
where $0<\alpha\leq 1$ and $\lambda>1$ capture diminishing sensitivity—concave curvature in the gain domain and a convex curvature in the loss domain—and loss aversion, respectively. Here, $x$ denotes the rank of lifter $i$ at the end of the stage and $r$ is the intermediate rank of $i$ before this stage. In our application, the change in rank $(x-r)$ takes values in $\{-1,0,1\}$.

### v4

@@ 2.1 Data source
We use data from OpenPowerlifting (2024), a community service project that creates a permanent and open archive of world powerlifting data. [^5] https://gitlab.com/openpowerlifting. Data accessed on July 1, 2023. The OpenPowerlifting database includes a wide range of fields from official powerlifting competitions. It contains competition details (federation name, date and location, and level: local, national, or international), athlete information (name, age, weight class, and gender), lift results for squat, bench press, and deadlift (attempted weights and success/failure status), Wilks score and other relative strength metrics, and rankings and historical records within competitions.
In this paper, we focus on bench press competitions, which represent the most popular division and are less complex than Squat-Bench-Deadlift (SBD) composite competitions. We use all official attempt data for lifters aged 15-69 in competition categories based on age class, weight class, and gender, where the number of participants is at least two to capture competitive pressure.
@@
Sources: The OpenPowerlifting database.
@@
Personal best weights of lifters who have never participated in an official bench press competition are recorded as zero. In the empirical analysis, we focus on the second and third attempts and treat the realized outcome in the first attempt as the personal best for such lifters because the success probability in the first attempt is 86% for raw lifters, 77% for single-ply lifters, and 70% for multi-ply lifters. [^6] Appendix A.2 shows the supplemental analysis of the first attempts, which shows similar findings to our main results on the second and third attempts.
@@ 2.2 Game setting
The objective of each lifter is to lift the maximum weight possible from a prone position on a flat bench to improve their own rank. In case of a tie (same weight lifted), the lighter lifter or the lifter who achieved the weight first (depending on the federation’s rules) is ranked higher. Competitors are divided into categories based on weight class, age class, equipment class, and gender to ensure fair competition. [^7] See details in the International Powerlifting Federation (IPF) Technical Rules Book.
@@
Each lifter has three attempts. [^8] Lifting Procedure for each lifter at each attempt is officially divided into the following. 1. Setup: The lifter lies on their back on the bench, with feet flat on the floor or on the bench’s footrests, ensuring contact with the bench throughout the lift. 2. Grip: The lifter grips the barbell, usually slightly wider than shoulder-width apart. 3. Unracking: With the help of spotters if needed, the lifter unracks the barbell and holds it with arms fully extended above their chest. 4. Lowering: The barbell is lowered in a controlled manner to touch the chest, ensuring it is stable and paused for a moment. 5. Press Command: The judge gives a “Press” command once the barbell is motionless on the chest. The lifter then presses the barbell back to the starting position with arms fully extended. 6. Rack Command: Upon completion, and with the barbell in a stable position, the judge gives the “Rack” command, signaling the lifter to return the barbell to the rack. Each attempt is judged based on control, stability, and completion of lifting. A lift is considered successful if the lifter adheres to all rules and completes the lift according to the judge’s commands, as determined by at least two of the three referees. After completing an attempt, the lifter must declare the weight for their next attempt within a specific time frame (usually 1 minute after their attempt). Once a weight is declared, it generally cannot be decreased, only increased. [^9] Lifters therefore tend to begin with relatively conservative weights and choose heavier weights in later attempts. This pattern can be interpreted as a natural strategic response to the need to avoid finishing with no recorded result, implying greater risk-taking in later rounds. The order of lifters is adjusted after each round of attempts based on the declared weights for the next attempt, again proceeding from lightest to heaviest. This ensures that the competition is efficient and fair.
@@
Notes: The box indicates that the lifter appears on the stage. $W_{it}^{j}$ is lifter $i$’s attempt weight in the $j$-th attempt in competition $t$.
To highlight the information available to each lifter at each attempt, consider a competition with three lifters (A, B, and C). The timeline of the first and second attempts is shown in Figure 1. For the opening attempts, lifters $A$, $B$, and $C$ declare 100 kg, 105 kg, and 110 kg before the competition. The order is $A$, $B$, and $C$. Lifter $A$ attempts 100 kg, and the outcome is realized. Next, lifter $B$ attempts 105 kg. At the same time, lifter $A$ declares the next attempt weight (110 kg) within one minute, conditional on $A$’s own first outcome and on $B$’s and $C$’s attempt weights, before $B$’s outcome is realized because lifter $B$ is attempting 105 kg at that time. Lifter $B$’s outcome is realized. Next, lifter $C$ attempts 110 kg. At the same time, lifter $B$ declares the next attempt weight (115 kg) within one minute, knowing $B$’s own first outcome, $A$’s outcome and next attempt, and $C$’s attempt weight, before $C$’s outcome is realized because lifter $C$ is attempting 110 kg at that time. Lifter $C$’s outcome is realized. Lifter $C$ declares the next attempt weight (112.5 kg) within one minute, knowing $A$’s and $B$’s outcomes, their next attempts, and $C$’s own first outcome. For the second attempts, the new order is $A$ (110 kg), $C$ (112.5 kg), and $B$ (115 kg). After the same procedure, the order for the third attempt is similarly determined by the declared weights in the second attempt.
2.3 Reference point in bench press competition
A wide range of reference points have been discussed in the literature, including the status quo, past experience, focal outcomes, aspirations, expectations, and norms (O’Donoghue and Sprenger 2018). In bench press competitions, potential candidates include personal best records, salient numbers, and opponents’ outcomes. In this study, however, we focus on a lifter’s current ranking within the competition as the reference point. Under this definition, competitive outcomes are naturally evaluated in terms of whether a lifter improves upon, maintains, or loses their current rank as the competition unfolds.
Bench press competitions are explicitly rank-based. Athletes compete to achieve a higher placement, and rankings are determined by the outcomes of a finite sequence of attempts. Because each competition consists of three attempts, intermediate rankings become observable from the second attempt onward. This allows lifters to assess how their relative position may change in response to rivals’ declared weights and realized outcomes, making current ranking a particularly salient benchmark for decision-making during the competition.
The sequential structure of the competition plays a crucial role in this setting. As discussed in Section 2.4, the timing of information revelation regarding rivals’ attempts and outcomes generates exogenous variation in competitive pressure relative to the current ranking. While individual lifters may differ in the extent to which they attend to various potential reference points, focusing on current ranking provides a natural and empirically tractable way to study how pressure arising from relative position affects risk-taking behavior and performance in a sequential competition.
@@ 2.4 Exogenous pressure
The above sequential game structure, in which each lifter does not know subsequent attempts and outcomes, provides exogenous variation in both attempt-weight choice and lifting success, analogous to lottery choice and realized payoff. We focus on the moves of the closest rivals to each lifter, namely the immediately lower- and higher-ranked rivals, which are critical for determining the final rank in the competition.
@@
For second-attempt weight choice, each lifter observes the first-attempt outcome and the second-attempt weight of the immediately lower-ranked rival. These are uncorrelated with the lifter’s unobservable characteristics, such as body condition on the competition date, because the rival declares the second-attempt weight before observing the outcome of the focal lifter’s second attempt, which—if observable—could serve as a proxy for the focal lifter’s condition. The exogenous pressure affects the risk of being overtaken by the lower-ranked rival, that is, the risk of loss. Second, each lifter forms an expectation about the second-attempt weight of the immediately higher-ranked rival, although this value is not yet observed. We assume that each lifter can predict a higher-ranked rival’s action based on information about the rival’s past attempts and a rich set of covariates. [^10] The assumption is supported by our data in Section 7.2. Strategic reasoning in intermediate weight selections is also inherently limited, as each subsequent attempt must exceed the current attempt. While incorporating dynamic and strategic considerations is theoretically appealing, doing so introduces significant complexity and falls beyond the scope of our empirical analysis. Because these predictions are based solely on observable characteristics, they are independent of the lifter’s own unobserved factors, such as temporary performance conditions. Because the first-attempt weight of the higher-ranked rival is heavier than the focal lifter’s current best weight, overtaking that rival is more difficult. The exogenous pressure affects the challenge of overtaking the higher-ranked rival, that is, the risk of gain.
@@
During the second-attempt lifting stage, each lifter knows the first- and second-attempt outcomes of the immediately lower-ranked rival. These are uncorrelated with the lifter’s unobserved characteristics, such as body condition on the competition date, because the lower-ranked rival completes both attempts without observing the focal lifter’s second outcome. The exogenous pressure affects the risk of being overtaken by the lower-ranked rival, that is, the risk of loss. If the focal lifter fails the second attempt, the rival overtakes him. Second, each lifter also knows the second-attempt weight of the immediately higher-ranked rival. This declared weight is uncorrelated with the lifter’s unobserved characteristics because, at the time of declaration, the higher-ranked rival has not yet observed the focal lifter’s second outcome. If the focal lifter succeeds in the second attempt, he has a greater chance to overtake the rival. The exogenous pressure therefore affects the challenge of overtaking the higher-ranked rival, that is, the risk of gain. Note that this pressure matters only at the lifting stage.
@@ 2.5 Data pattern
We illustrate data patterns in attempt-weight choice and the success probability of an attempt conditional on the exogenous pressure defined above. For illustration, we show the most popular competition category (Male, Raw, 24-39 age class). Similar patterns appear in other competition categories. We omit the first-attempt patterns because the success probability at that stage is above 75%.
@@
In summary, the data suggest that when lifters face pressure in critical situations—such as attempting to improve their rank—they tend to choose more challenging weights. However, unlike the second attempt, where pressured lifters often maintain higher success rates, the third attempt shows a more substantial decline in overall success probabilities, suggesting increased difficulty in securing successful lifts under pressure.
@@
We examine the pressure exerted by rivals during competition. A natural question arises: who constitutes a rival for each lifter within a competition? Table 3 illustrates the distribution of rank changes for lifters between their second and third attempts. The majority of lifters maintain their rank, as shown by the high proportion of zero rank changes (67% in the second attempt and 72% in the third attempt). A smaller percentage of lifters experience a rank increase (+1+1) between attempts (14% in the second and 10% in the third), while the proportion of those gaining two or more ranks is notably lower (5% in the second and 3% in the third). Similarly, lifters dropping one rank (1-1) account for 8% in the second attempt and 9% in the third, with even fewer dropping two or more ranks (5% and 6%, respectively). Based on this distribution, we focus on the immediate competitors, specifically the lifter ranked directly below and the one ranked directly above, as these rivals most significantly influence rank changes.
@@
3 Conceptual framework
In this section, we develop a conceptual framework based on expected utility theory and prospect theory. Bench press competitions are organized in successive rounds, with interim rankings determined after each round. This setting naturally motivates a reference-dependent evaluation of competitive outcomes, where the interim rank serves as the reference point.
Following Tversky and Kahneman (1992), we adopt a rank-dependent value function $v(\cdot)$ with reference point $r$ defined as
$v(x)=\begin{cases}(x-r)^{\alpha}&\text{if }x\geq r,\\ -\lambda(r-x)^{\alpha}&\text{if }x<r,\end{cases}$
where $0<\alpha\leq 1$ and $\lambda>1$ capture diminishing sensitivity—concave curvature in the gain domain and convex curvature in the loss domain—and loss aversion, respectively. Here, $x$ denotes the rank of lifter $i$ at the end of a stage, and $r$ denotes lifter $i$’s interim rank prior to that stage. We adopt the convention that higher values correspond to better positions; accordingly, we reverse the ordering of the ordinal rank so that the best position takes the largest value. Under this convention, xr=+1x-r=+1 corresponds to an improvement by one position and xr=1x-r=-1 to being overtaken by one position. For simplicity, we focus on local competition and assume $x-r\in\{-1,0,1\}$. [^11] Table 3 shows that rank changes of two or more positions are uncommon (approximately 5% of observations), so this restriction captures the empirically relevant margin. Moreover, when $x-r\in\{-1,0,1\}$, we have $|x-r|^{\alpha}=|x-r|$ for any $0<\alpha\leq 1$, implying that local asymmetry is governed by $\lambda$. To maintain tractability and focus on empirically testable implications, we model the attempt and lifting stages separately and treat rivals’ decisions as given from the focal lifter’s perspective.  [^12] A fully dynamic equilibrium model with endogenous effort and strategic interaction would substantially complicate the framework and is beyond the scope of the present study.
3.1 Attempt stage
At the attempt stage, lifter $i$ chooses an attempt weight $w_{i}$. We focus on local competition among three lifters: lifter $i$, the lower-ranked rival $L$ immediately below $i$, and the higher-ranked rival $H$ immediately above $i$ in the interim ranking at the start of the round. Let $y_{i}$ denote lifter $i$’s highest successfully lifted weight prior to the current round. Since attempt weights cannot be reduced once declared, the feasible set satisfies $w_{i}\geq y_{i}$. For any declared attempt weight $w$, let $q_{i}(w)\in[0,1]$ denote lifter $i$’s probability of success. We assume that heavier weights are more difficult to lift, so that $q_{i}^{\prime}(w)<0$ for any $i$.
We first consider pressure from the lower-ranked rival. Suppose that the lower rival declares an attempt weight $w_{L}>y_{i}$. If lifter $i$ chooses $w_{i}$ such that $w_{L}>w_{i}$, lifter $i$ is overtaken if and only if the lower rival succeeds. The expected utility is then given by $EU\mid_{w_{L}>w_{i}}=-\lambda q_{L}(w_{L})\leq 0$. If instead lifter $i$ chooses $w_{i}>w_{L}$, overtaking occurs only when lifter $i$ fails and the lower rival succeeds. The expected utility becomes $EU\mid_{w_{i}>w_{L}}=-\lambda\bigl(1-q_{i}(w_{i})\bigr)q_{L}(w_{L})\leq 0$.
Since $0<1-q_{i}(w_{i})<1$, we have
$EU\mid_{w_{i}>w_{L}}>EU\mid_{w_{L}>w_{i}}\quad\text{and}\quad\frac{\partial EU\mid_{w_{i}>w_{L}}}{\partial w_{i}}=\lambda q_{L}(w_{L})q_{i}^{\prime}(w_{i})\leq 0,$
which implies that lifter $i$ maximizes expected utility by choosing an attempt weight just above $w_{L}$.
When $w_{L}<y_{i}$, no rank change occurs regardless of the success or failure outcomes of either lifter. In this case, expected utility satisfies $EU(w_{i})\mid_{y_{i}>w_{L}}=0$, and lifter $i$ is indifferent over all feasible choices of $w_{i}$.  [^13] Note that this situation depends on the lower rival’s choice and cannot be induced by lifter $i$’s own declaration.
Hypothesis 1.
When the lower rival’s declared attempt weight exceeds lifter $i$’s current best, an increase in competitive pressure from below—measured by $w_{L}-y_{i}$—induces lifter $i$ to choose a heavier attempt weight, i.e., to engage in greater risk-taking.
We next turn to pressure from the higher-ranked rival. The higher rival’s declared attempt weight $w_{H}$ is not directly observed by lifter $i$ at the time of the attempt stage. However, as discussed in Section 7.2, lifter $i$ can form expectations about $w_{H}$ based on observable characteristics.
If lifter $i$ chooses $w_{i}<y_{H}$, no rank change occurs regardless of the success or failure of $i$’s outcome. In this case, expected utility satisfies $EU\mid_{y_{H}>w_{i}}=0$, and lifter $i$ is indifferent over all such choices of $w_{i}$.
By contrast, choosing $w_{i}>y_{H}$ creates a possibility of overtaking the higher-ranked rival. We therefore compare the expected utility from alternative choices with $w_{i}>y_{H}$, taking as given lifter $i$’s expectations about the higher rival’s attempt weight $w_{H}$. Given an expectation of $w_{H}$, lifter $i$ may declare either $w_{i}^{-}\in(y_{H},w_{H})$ or $w_{i}^{+}>w_{H}$. If $w_{H}>w_{i}^{-}>y_{H}$, lifter $i$ overtakes the higher rival only if lifter $i$ succeeds and the higher rival fails. Expected utility is therefore $EU\mid_{w_{H}>w_{i}^{-}>y_{H}}=q_{i}(w_{i}^{-})\,\bigl(1-q_{H}(w_{H})\bigr)\geq 0$. If $w_{i}^{+}>w_{H}\geq y_{H}$, lifter $i$ overtakes the higher rival whenever lifter $i$ succeeds. Expected utility is then $EU\mid_{w_{i}^{+}>w_{H}}=q_{i}(w_{i}^{+})\geq 0$.
Since choosing $w_{i}<y_{H}$ yields zero expected utility while $w_{i}>y_{H}$ yields weakly positive expected utility, lifter $i$ prefers to select an attempt weight exceeding $y_{H}$. Among choices with $w_{i}>y_{H}$, expected utility is decreasing in $w_{i}$ in both relevant regions. In particular,
$\frac{\partial EU\mid_{w_{i}^{+}>w_{H}}}{\partial w_{i}^{+}}=q_{i}^{\prime}(w_{i}^{+})<0\quad\text{and}\quad\frac{\partial EU\mid_{w_{H}>w_{i}^{-}}}{\partial w_{i}^{-}}=q_{i}^{\prime}(w_{i}^{-})\,\bigl(1-q_{H}(w_{H})\bigr)\leq 0.$
Therefore, the optimal choice lies at the boundary of each region: either just above $y_{H}$ or just above the expected $w_{H}$.
Comparing these two boundary choices, lifter $i$ prefers declaring a weight just above the expected $w_{H}$ rather than just above $y_{H}$ if and only if
$\frac{q_{i}(w_{i}^{+})}{q_{i}(w_{i}^{-})}>1-q_{H}(w_{H}).$
When this condition holds, lifter $i$ optimally declares an attempt weight slightly exceeding the expected $w_{H}$; otherwise, lifter $i$ declares a weight slightly exceeding $y_{H}$.
Hypothesis 2.
When the higher rival’s expected attempt weight exceeds lifter $i$’s current best, an increase in competitive pressure from above—measured by $w_{H}-y_{i}$—induces lifter $i$ to choose an attempt weight at least as heavy as $y_{H}$. Moreover, when the higher rival is expected to succeed with high probability, or when lifter $i$’s own success probability at $y_{H}$ is low, lifter $i$ is more likely to engage in even greater risk-taking by declaring an attempt weight exceeding the higher rival’s expected attempt.